### v1

# Getting it Right: Improving Spatial Consistency in Text-to-Image Models

Abstract
One of the key shortcomings in current text-to-image (T2I) models is their inability to consistently generate images which faithfully follow the spatial relationships specified in the text prompt. In this paper, we offer a comprehensive investigation of this limitation, while also developing datasets and methods that achieve state-of-the-art performance. First, we find that current vision-language datasets do not represent spatial relationships well enough; to alleviate this bottleneck, we create SPRIGHT, the first spatially-focused, large scale dataset, by re-captioning 6 million images from 4 widely used vision datasets. Through a 3-fold evaluation and analysis pipeline, we find that SPRIGHT largely improves upon existing datasets in capturing spatial relationships. To demonstrate its efficacy, we leverage only $\sim$0.25of SPRIGHT and achieve a 22% improvement in generating spatially accurate images while also improving the FID and CMMD scores. Secondly, we find that training on images containing a large number of objects results in substantial improvements in spatial consistency. Notably, we attain state-of-the-art on T2I-CompBench with a spatial score of 0.2133, by fine-tuning on <500 images. Finally, through a set of controlled experiments and ablations, we document multiple findings that we believe will enhance the understanding of factors that affect spatial consistency in text-to-image models. We publicly release our dataset and model to foster further research in this area.
Figure 1: We find that existing vision-language datasets do not capture spatial relationships well. To alleviate this shortcoming, we synthetically re-caption $\sim$6M images with a spatial focus, and create the SPRIGHT (SPatially RIGHT) dataset. Shown above are samples from the COCO Validation Set, where text in red denotes ground-truth captions and text in green are corresponding captions from SPRIGHT.
1 Introduction
The development of text-to-image (T2I) diffusion models such as Stable Diffusion [54] and DALL-E 3 [45] has led to the growth of image synthesis frameworks that are able to generate high resolution photo-realistic images. These models have been adopted widely in downstream applications such as video generation [61], image editing [21], robotics [16], and more. Multiple variations of T2I models have also been developed, which vary according to their text encoder [6], priors [53], and inference efficiency [41]. However, a common bottleneck that affects all of these methods is their inability to generate spatially consistent images: that is, given a natural language prompt that describes a spatial relationship, these models are unable to generate images that faithfully adhere to it.
In this paper, we present a holistic approach towards investigating and mitigating this shortcoming through diverse lenses. We develop datasets, efficient and effective training techniques, and explore multiple ablations and analyses to understand the behaviour of T2I models towards prompts that contain spatial relationships.
Our first finding reveals that existing vision-language (VL) datasets lack sufficient representation of spatial relationships. Although frequently used in the English lexicon, we find that spatial words are scarcely found within image-text pairs of the existing datasets. To alleviate this shortcoming, we create the “SPRIGHT” (SPatially RIGHT) dataset, the first spatially-focused large scale dataset. Specifically, we synthetically re-caption $\sim$6 million images sourced from 4 widely used datasets, with a spatial focus (Section 3). As shown in Figure 1, SPRIGHT captions describe the fine-grained relational and spatial characteristics of an image, whereas human-written ground truth captions fail to do so. Through a 3-fold comprehensive evaluation and analysis of the generated captions, we benchmark the quality of the generated captions and find that SPRIGHT largely improves over existing datasets in its ability to capture spatial relationships. Next, leveraging only $\sim$0.25of our dataset, we achieve a 22% improvement on the T2I-CompBench [23] Spatial Score, and a 31.04% and 29.72% improvement in the FID [22] and CMMD scores [24], respectively.
Our second finding reveals that significant performance improvements in spatial consistency of a T2I model can be achieved by fine-tuning on images that contain a large number of objects. We achieve state-of-the-art performance, and improve image fidelity, by fine-tuning on <500 image-caption pairs from SPRIGHT; training only on images that have a large number of objects. As investigated in VISOR [20], models often fail to generate the mentioned objects in a spatial prompt; we posit that by optimizing the model over images which have a large number of objects (and consequently, spatial relationships), we teach it to generate a large number of objects, which positively impacts its spatial consistency. In addition to improving spatial consistency, our model achieves large gains in performance across all aspects of T2I generation; generating correct number of distinct objects, attribute binding and accurate generation in response to complex prompts.
We further demonstrate the impact of SPRIGHT by benchmarking the trade-offs achieved with long and short spatial captions, as well spatially focused and general captions. Finally, we take the first steps towards discovering layer-wise activation patterns associated with spatial relationships, by examining the representation space of CLIP [50] as a text encoder. To summarize, our contributions are as follows:
  1. 1.
    We create SPRIGHT, the first spatially focused, large scale vision-language dataset by re-captioning $\sim$6 million images from 4 widely used existing datasets.To demonstrate the efficacy of SPRIGHT, we fine-tune baseline Stable Diffusion models on a small subset of our data and achieve performance gains across multiple spatial reasoning benchmarks while improving the corresponding FID and CMMD scores.
  2. 2.
    We achieve state-of-the-art performance on spatial relationships by developing an efficient training methodology; specifically, we optimize over a small number (<500) of images which consists of a large number of objects, and achieve a 41% improvement over our baseline model.
  3. 3.
    Through multiple ablations and analyses, we present our diverse findings related to spatial relationships: the impact of long captions, the trade-off between spatial and general captions, layer-wise activations of the CLIP text encoder, effect of training with negations and improvements over attention maps.
2 Related Work
Text-to-image generative models. Since the initial release of Stable Diffusion [54] and DALL-E [52], different classes of T2I models have been developed, all optimized to generate highly realistic images corresponding to complex natural prompts. Models such as PixArt-Alpha [6], Imagen [55], and ParaDiffusion [62] move away from the CLIP text encoder, and explore traditional language models such as T5 [51] and LLaMA [59] to process text prompts. unCLIP [53] based models have led to multiple methods [48, 33] that leverage a CLIP-based prior as part of their diffusion pipeline. Recent works have also explored improvements in model compression [35, 27] and inference efficiency [41, 42].
Spatial relationships in T2I models. Benchmarking the failures of T2I models on spatial relationships has been well explored by VISOR [20], T2I-CompBench [23], GenEval [17], and DALL-E Eval [9]. Both training-based and test-time adaptations have been developed to specifically improve upon these benchmarks. Control-GPT [68] finetunes a ControlNet [67] model by generating TikZ code representations with GPT-4 and optimizing over grounding tokens to generate images. SpaText [1], GLIGEN[34], and ReCo [64] are additional training-based methods that introduce additional conditioning in their fine-tuning process to achieve better spatial control for image generation. LLM-Grounded Diffusion [36] is a test-time multi-step method that improves over layout generated LLMs in an iterative manner. Layout Guidance [7] restricts objects to their annotated bounding box locations through refinement of attention maps during inference. LayoutGPT [15] creates initial layout guidance from LLMs in the form of CSS code, and then uses layout-to-image models to create indoor scenes.
Synthetic captions for T2I models. The efficacy of using descriptive and detailed captions has recently been explored by DALL-E 3 [45], PixArt-Alpha [6] and RECAP [58]. DALL-E 3 builds an image captioning module by jointly optimizing over a CLIP and language modeling objective. RECAP fine-tunes an image captioning model (PALI [8]) and reports the advantages of fine-tuning the Stable Diffusion family of models on long, synthetic captions. PixArt-Alpha also re-captions images from the LAION [56] and Segment Anything [28] datasets; however their key focus is to develop descriptive image captions. On the contrary, our goal is to develop captions that explicitly capture the spatial relationships seen in the image.
Dataset % of Spatial Phrases
left right above below front behind next close far small large
COCO 0.16 0.47 0.61 0.15 3.39 1.09 6.17 1.39 0.19 3.28 4.15
   + SPRIGHT 26.80 23.48 21.25 5.93 41.68 21.13 36.98 15.85 1.34 48.55 61.80
CC-12M 0.61 1.45 0.40 0.19 1.40 0.43 0.54 0.94 1.07 1.44 1.44
   + SPRIGHT 24.53 22.36 20.42 6.48 41.23 14.37 22.59 12.9 1.10 43.49 66.74
LAION 0.27 0.75 0.16 0.05 0.83 0.11 0.24 0.67 0.91 1.03 1.01
   + SPRIGHT 24.36 21.70 14.27 4.07 42.92 16.38 26.93 13.05 1.16 49.59 70.27
Segment Anything 0.02 0.07 0.27 0.06 5.79 0.19 3.24 7.51 0.05 0.85 10.58
   + SPRIGHT 18.48 15.09 23.75 6.56 43.5 13.58 33.02 11.9 1.25 52.19 80.22
Table 1: Compared to ground truth annotations, SPRIGHT consistently improves the presence of relational and spatial relationships captured in its captions, across diverse images from different datasets.
3 The SPRIGHT Dataset
We find that current vision-language (VL) datasets do not contain “enough” relational and spatial relationships. Despite being frequently used in the English vocabulary  [^1] https://www.oxfordlearnersdictionaries.com/us/wordlists/oxford3000-5000, words like “left/right”, “above/behind” are scarce in existing VL datasets. This holds for both annotator-provided captions, e.g., COCO [37], and web-scraped alt-text captions, e.g., LAION [57]. We posit that the absence of such phrases is one of the fundamental reasons for the lack of spatial consistency in current text-to-image models. Furthermore, language guidance is now being used to perform mid-level [66, 63] and low-level [70, 29] computer vision tasks. This motivates us to create the SPRIGHT (SPatially RIGHT) dataset, which explicitly encodes fine-grained relational and spatial information found in images.
3.1 Creating the SPRIGHT Dataset
We re-caption $\sim$6M images from 4 well-known VL datasets:
  • CC-12M [4] : We re-caption a total of 2.3 million images from the CC-12M dataset, filtering out images of resolution less than 768.
  • Segment Anything (SA) [28] : We select Segment Anything as we find that most images in it encapsulates a large number of objects; i.e. larger number of spatial relationships can be captured from a given image. We re-caption 3.5 million images as part of our re-captioning process. Since SA does not have ground-truth captions, we also generate its general captions using the CoCa [65] model.
  • COCO [37] : We re-caption images from the COCO Validation Set, totalling to $\sim$ 40,000 images.
  • LAION-Aesthetics [^2] https://laion.ai/blog/laion-aesthetics/ : We consider 50,000 images from the LAION-Aesthetics dataset and do not include more images as the dataset is currently publicly unavailable. (We also do not release the upsampled captions at this time, for the same reason.)
We leverage LLaVA-1.5-13B [38] to produce synthetic spatial captions, and use the following prompt to create the SPRIGHT Dataset:
Using 2 sentences, describe the spatial relationships seen in the image. You can use words like left/right, above/below, front/behind, far/near/adjacent, inside/outside. Also describe relative sizes of objects seen in the image.
We also explore other relevant open models such as FUYU-8B [2] and IDEFICS-9B-Instruct [32], but find LLaVA to be the most consistent at generating spatially accurate captions.
3.2 Impact of SPRIGHT
As shown in Table 1, we find that SPRIGHT greatly enhances the presence of spatial phrases across all relationship types, on all the datasets. Specifically, considering 11 spatial and relational relationships, ground-truth captions of COCO and LAION only capture 21.05% and 6.03% of relationships; on the other hand, SPRIGHT captures 304.79% and 284.7%, respectively. This shows that captions in VL datasets largely lack the presence of spatial relationships, and that SPRIGHT is able to improve upon this shortcoming by almost always capturing spatial relationships in every sentence.
Our captions offer several improvements beyond the spatial aspects: (i) As depicted in Table 2 we improve the overall linguistic quality compared to the original captions, and (ii) we identify more objects and amplify their occurrences as illustrated in Figure 2; where we plot the top 10 objects present in the original COCO Captions and find that we significantly upsample their corresponding presence in SPRIGHT.
Dataset Average / caption
Nouns Adjective Verbs Tokens
COCO 3.00 0.83 0.04 11.28
   + SPRIGHT 14.31 3.82 0.15 68.22
CC-12M 3.35 1.36 0.26 22.93
   + SPRIGHT 13.99 4.36 0.16 67.41
LAION 1.78 0.70 0.11 12.49
   + SPRIGHT 14.32 4.53 0.14 69.74
Segment Anything 3.1 0.79 0.01 9.88
   + SPRIGHT 13.42 4.65 0.12 63.9
Table 2: In addition to improving the presence of spatial relationships, SPRIGHT also enhances the overall linguistic variations and diversity of captions in comparison to their original versions.
Figure 2: (Left) On comparing word cloud representations on the COCO dataset, we find that SPRIGHT significantly amplifies the presence of spatial relationships. (Right) Correspondingly, we also improve fine-grained object identification and their presence in our captions.
Figure 3: The distribution of ratings assigned to 100 captions from SPRIGHT by GPT-4(V), on a scale from 1 to 10.
3.3 Dataset Validation
We perform 3 levels of evaluation to validate the correctness of the captions present in SPRIGHT:
Category # Examples Accuracy (%)
Overall 88.9
FAITHScore
Entities 149,393 91.4
Relations 167,786 85.8
Colors 10,386 83.1
Counting 59,118 94.5
Other 29,661 89.0
Spatial 45,663 83.6
Table 3: FAITHScore caption evaluation of our SPRIGHT dataset. On a sample of 40,000 random captions, SPRIGHT obtains an 88.9% accuracy, comparable with the reported 86% and 94% on LLaVA-1k and MSCOCO-Captions, respectively [25]. On the subset of atomic claims about spatial relations, SPRIGHT is correct 83.6% of the time.
1. FAITHScore. Following [25], we leverage a large language model to deconstruct generated captions into atomic (simple) claims that can be individually and independently verified in a Visual Question Answering (VQA) format. We randomly sample 40,000 image-generated caption pairs from our dataset, and prompt GPT-3.5-Turbo to identify descriptive phrases (as opposed to subjective analysis that cannot be verified from the image) and decompose the descriptions into atomic statements. These atomic statements are then passed to LLaVA-1.5-13B for verification, and correctness is aggregated over 5 categories: entity, relation, colors, counting, and other attributes. We also measure correctness on spatial-related atomic statements, i.e., those containing one of the keywords left/right, above/below, near/far, large/small and background/foreground. As shown in Table 3 the captions are on average 88.9% correct, with spatially-focused relations, being 83.6% correct. Since there is some uncertainty about bias induced by using LLaVA to evaluate LLaVA-generated captions, we also verify the caption quality in other ways, as described next.
2. GPT-4 (V). Inspired by recent methods [45, 71], we perform a small-scale study on 100 images to evaluate our captions with GPT-4(V) [46]. We prompt GPT-4(V) to rate each caption between a score of 1 to 10, especially focusing on the correctness of the spatial relationships captured. We present the distribution of scores in Figure 3 – the mean and median rating is 6.41 and 7.00 respectively. We present the GPT-4 (V) prompt used in Appendix A.
3. Human Annotation. We also annotate a total of 3,000 images through a crowd-sourced human study, where each participant annotates a maximum of 30 image-text pairs. As evidenced by the average number of tokens in Table 1, most captions in SPRIGHT have >1 sentences. Therefore, for fine-grained evaluation, we randomly select 1 sentence, from a caption in SPRIGHT, and evaluate its correctness for a given image. Across 149 responses, we find the metrics to be: correct=1840 and incorrect=928, yielding an accuracy of 66.57%.
4 Improving Spatial Consistency
In this section, we leverage SPRIGHT in an effective and efficient manner, and describe methodologies that significantly advance spatial reasoning in T2I models. Our experimental setting is as follows: Our base model is Stable Diffusion v2.1  [^3] https://huggingface.co/stabilityai/stable-diffusion-2-1; our training and validation set consists of 13,500 and 1,500 images respectively, randomly sampled in a 50:50 split between LAION-Aesthetics and Segment Anything. Each image is paired with both, a general and a spatial caption (from SPRIGHT). During fine-tuning, for each image, we randomly choose one of the given caption types in a 50:50 ratio. We fine-tune the U-Net and the CLIP Text Encoder as part of our training, both with a learning rate $5\times 10^{-6}$ optimized by AdamW [40] and a global batch size of 128. While we train the U-Net for 15,000 steps, the CLIP text encoder remains frozen during the first 10,000 steps. We develop our code-base on top of the Diffusers library [60].
[…]
uncond cond 1 2 3 4
SD 2.1 47.83 30.25 63.24 64.42 35.74 16.13 4.70 0.1507 21.646 0.703
   + SPRIGHT 53.59 36.00 67.16 66.09 44.02 24.15 9.13 0.1840 14.925 0.494
Table 4: Quantitative metrics across multiple spatial reasoning and image fidelity metrics, demonstrating the effectiveness of high quality spatially-focused captions in SPRIGHT. Green indicates results of the model fine-tuned on SPRIGHT. To calculate FID, we use cfg = 3.0 and 7.0 for the baseline and the fine-tuned model, respectively
Figure 4: Illustative examples from our model, as described in Section 4.1, on evaluation prompts from the T2I-CompBench benchmark.
Figure 5: Generated images from our model, as described in Section 4.1, on prompts which contain multiple objects and complex spatial relationships. We curate these prompts from ChatGPT.
4.1 Improving upon Baseline Methods
We present our quantitative results on the spatial relationship benchmarks (VISOR [20], T2I-CompBench [23]) and image fidelity metrics in Table 4. To account for the inconsistencies associated with FID [10, 47], we also report results on CMMD [24]. Across all metrics, we significantly improve upon the base model by fine-tuning on <15k images. We conclude that the dense, spatially focused captions in SPRIGHT provide effective spatial guidance to T2I models, and alleviate the need to scale up fine-tuning on a large number of images. Illustrative examples on evaluation prompts from T2I-CompBench are shown in Figure 4, while images corresponding to complex prompts are presented in Figure 5. As shown in Figure 5, we find that the model captures complex spatial relationships (top right), relative sizes (large) and patterns (swirling).
4.2 Efficient Training Methodology
We devise an additional efficient training methodology, which achieves state-of-the-art performance on the spatial aspect of the T2I-CompBench Benchmark. We hypothesize that (a) images that capture a large number of objects inherently also contain multiple spatial relationships; and (b) training on these kinds of images will optimize the model to consistently generate a large number of objects, given a prompt containing spatial relationships; a well-documented failure mode of current T2I models [20].
We partition our dataset of <15k images, as described in Section 4, and bucket them into multiple splits, where each split has images constrained with the # of objects present in them. We achieve this by leveraging the open-world image tagging model Recognize Anything [69]. Subsequently, we train respective models using training images solely from a given split, and benchmark their performance. We keep the same hyper-parameters as before, only initiating training of the CLIP Text Encoder from the beginning. For our dataset split of <15k images, we find the median # of objects/ image to be 11. Thus, we create 5 splits, train corresponding models and benchmark them in Table 5. We find that with the increase in the # of objects / image, we are iteratively able to improve upon the spatial fidelity.
[…]
<6 444 0.1309
<11 1346 0.1468
11 1346 0.1667
>11 1346 0.1613
> 18 444 0.2133
Table 5: Across all reported methods, we achieve state-of-the-art performance on the T2I-CompBench Spatial Score. This is achieved by fine-tuning SD 2.1 on 444 image-caption pairs from the SPRIGHT dataset; where each image has >18 objects.
[…]
uncond cond 1 2 3 4
SD 2.1 47.83 30.25 63.24 64.42 35.74 16.13 4.70 0.1507 21.646 0.703
[…]
Table 6: Comparing baseline SD 2.1 with our state-of-the-art model, across multiple spatial reasoning and image fidelity metrics, as described in Section 4.2 Green indicates results from our model. To calculate FID, we use cfg = 3.0 and 7.5 for baseline and our model, respectively
Method OA (%) VISOR (%)
uncond cond 1 2 3 4
GLIDE [44] 3.36 1.98 59.06 6.72 1.02 0.17 0.03
GLIDE + CDM [39] 10.17 6.43 63.21 20.07 4.69 0.83 0.11
CogView2 [12] 18.47 12.17 65.89 33.47 11.43 3.22 0.57
DALLE-mini [11] 27.10 16.17 59.67 38.31 17.50 6.89 1.96
DALLE-2 [53] 63.93 37.89 59.27 73.59 47.23 23.26 7.49
Structured Diffusion [14] 28.65 17.87 62.36 44.70 18.73 6.57 1.46
Attend-and-Excite [5] 42.07 25.75 61.21 49.29 19.33 4.56 0.08
Ours (<500 images) 60.68 43.23 71.24 71.78 51.88 33.09 16.15
Table 7: Results on the VISOR Benchmark. Our model outperforms existing methods, on all aspects related to spatial relationships, consistently generating spatially accurate images as shown by the high VISOR [1-4] values.
[…]
CLIP retrieval [3] 0.35 0.89 0.22 0.37 0.62 0.03 0.00
minDALL-E [31] 0.23 0.73 0.11 0.12 0.37 0.02 0.01
SD 1.5 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD 2.1 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL [49] 0.55 0.98 0.74 0.39 0.85 0.15 0.23
PixArt-Alpha [6] 0.48 0.98 0.50 0.44 0.80 0.08 0.07
Ours (<500 images) 0.51 0.99 0.59 0.49 0.85 0.11 0.15
Table 8: Results on the GenEval Benchmark. In addition to improving spatial relationships, we also alleviate performance while generating 1/2 objects along with generating the correct number of objects.
Our major finding is that, with 444 training images and spatial captions from SPRIGHT, we are able to attain a 41% improvement over baseline SD 2.1 and achieve state-of-the art across all reported models on the T2I-CompBench spatial score. In Table 6, we find that compared to SD 2.1, we largely improve on all aspects of the VISOR score, while improving the ZS-FID and CMMD score on COCO-30K images by 25.39% and 27.16%, respectively. On VISOR, our key findings are : (a) We increase the Object Accuracy (OA) score by 26.86%, indicating that we are much better at generating objects mentioned in the input prompt, (b) Our VISOR4 score of 16.15% denotes that for a given input prompt, we consistently generate a spatially accurate image (Figure 6). Our model outperforms existing methods on the VISOR benchmark (Table 7), with a notable improvement of 115.61% improvement on the VISOR4, denoting superior consistency in generating spatially correct images.
Next, we compare our model’s performance on the GenEval [17] benchmark (Table 8), and find that in addition to improving spatial relationship (see Position), we also enhance the ability to generate 1 and 2 objects, along with generating the correct number of objects. This further emphasizes our model’s ability to correctly generate objects mentioned in the input prompt. We present additional results on T2I-CompBench in Appendix B.
Figure 6: Generated images from our model, as described in Section 4.2, on evaluation prompts from T2I-CompBench. We find that for a given text prompt, our model consistently generates spatially accurate images.
Throughout our experiments, we find that our training approach not only preserves but also enhances the non-spatial aspects associated with a text-to-image model. We present illustrative images on VISOR prompts in Appendix C.
5 Ablations and Analyses
To fully ascertain the impact of spatially-focused captions in SPRIGHT, we experiment with multiple nuances of our dataset and the corresponding T2I pipeline. Unless stated otherwise, the experimental setup is the same as Section 4.
Figure 7: Illustrative comparisons between models trained on varying ratio of spatial experiments. We find that when trained on 50% and 75% ratio of spatial captions, model achieve their optimal performance. For each prompt, the images are sampled from the same seed, for a fair comparison.
5.1 Optimal Ratio of Spatial Captions
To understand the impact of spatially focused captions in comparison to ground-truth (non-spatial) captions, we fine-tune different models by varying the % of spatial captions as (25/50/75/100). Our results suggest that the model trained on 50% spatial captions achieves the most optimal scores on T2I-CompBench (Table 9). We find that models trained on only 25% of spatial captions suffer largely from incorrect spatial relationships whereas the model trained only on spatial captions fails to generate both the mentioned objects in the input prompt. Figure 7 shows illustrative examples.
[…]
25 0.154
50 0.178
75 0.161
100 0.140
Table 9: Comparing the T2I-CompBench Spatial Scores between models trained on varying ratio of spatial captions. Fine-tuning on a ratio of 50% and 75% of spatial captions yield the most optimal results on spatial relationships.
5.2 Impact of Long and Short Spatial Captions
We also compare the effect of fine-tuning with shorter and longer variants of spatial captions. We create the shorter variants by randomly sampling 1 sentence from the longer caption, and fine-tune multiple models, with different setups. Across, all setups, we find (Table 10) that longer captions perform better than their shorter counterparts. In fact, we find that CLIP fine-tuning hurts performance while using shorter captions, but has a positive impact on longer captions. This potentially happens because fine-tuning CLIP enables T2I models to generalize better to longer captions, which are out-of-distribution at the onset of training by virtue of them being initially pre-trained on short(er) captions from datasets such as LAION.
[…]
Long Captions Short Captions
SD 1.5, w/o CLIP FT 0.0910 0.0708
SD 2.1, w/o CLIP FT 0.1605 0.1420
SD 2.1, w/ CLIP FT 0.1777 0.1230
Table 10: Comparing the T2I-CompBench Spatial Scores between models trained on long and short spatial captions. Across multiple fine-tuning setups, we find that longer spatial captions lead to better improvements in spatial consistency.
5.3 Investigating the CLIP Text Encoder
The CLIP Text Encoder enables semantic understanding of the input text prompts in the Stable Diffusion model. As we fine-tune CLIP on the spatial captions from SPRIGHT, we investigate the various nuances associated with it:
Figure 8: Comparing layer-wise learned representations between the Baseline CLIP (X-axis) and its fine-tuned version on SPRIGHT (Y-axis). When compared on spatial captions, we find that distinct representations are learned in the output attention projections and the MLP layers, while layer norm layers higher levels of similarities. The gap in the learned representations increases when complex prompts are used, indicating that longer spatial prompts in SPRIGHT indeed lead to diverse representations in the CLIP embedding space.
Centered Kernel Alignment: The CKA metric [43, 30] allows us to compare the different layer-wise representations learned by two neural networks. Figure 8 illustrates different representations learned by baseline CLIP, compared against the one trained on SPRIGHT. We compare layer activations across 50 simple and complex prompts and aggregate representations from all the layers. Our findings reveal that the MLP and output attention projection layers play a larger role in enhancing spatial comprehension, as opposed to layers such as the layer norm. This distinction becomes more pronounced with complex prompts, emphasizing that the longer prompts from SPRIGHT lead to more diverse embeddings being learned within the CLIP space.
[…]
above 0.9225 0.8674
below 0.9259 0.8673
to the left of 0.9229 0.8658
to the right of 0.9223 0.8528
in front of 0.9231 0.8417
behind 0.9289 0.8713
Table 11: We find that CLIP fine-tuned on SPRIGHT is able to differentiate the spatial nuances present in a textual prompt. While Baseline CLIP shows a high similarity for spatially different prompts, SPRIGHT enables better fine-grained understanding.
Improving Semantic Understanding
: To evaluate semantic understanding of the fine-tuned CLIP, we conduct the following experiment: Given an original prompt containing a spatial phrase and 2 objects, we modify the prompt by switching the objects. For example, an original prompt - "an airplane above an apple" is modified to "an apple above an airplane". Our rationale is that, although both sentences are semantically close to each other, they are spatially distinct. we posit that a better model should be able to discern the fine-grained spatial distinction that are conveyed through both the sentences. We perform this experiment for 6 kinds of spatial relationships, for a total of $\sim$ 37k sentences. For each model, we compute the cosine similarity between the pooled layer outputs of the original and modified prompts. We find, as shown in Table 11, that fine-tuned CLIP indeed is able to differentiate better between the spatial prompts. And, it does not suffer from catastrophic forgetting, which would have yielded a very low cosine similarity.
Figure 9: Visualising the cross-attention relevancy maps for baseline (top row) and fine-tuned model (bottom row) on SPRIGHT. Images in red are from baseline model while images in green are from our model.
5.4 Training with Negation
Dealing with negation remains a challenge for multimodal models as reported by previous findings on Visual Question Answering and Reasoning [18, 19, 13]. Thus, in this section, we investigate the ability of T2I models to reason over spatial relationships and negations, simultaneously. Specifically, we study the impact of training a model with ‘‘A man is not to the left of a dog’’ as a substitute to ‘‘A man is to the right of a dog’’. To create such captions, we post-process our generated captions and randomly replace spatial occurrences with their negation counter-parts, and ensure that the semantic meaning of the sentence remains unchanged. Training on such a model, we find slight improvements in the spatial score, both while evaluating on prompts containing only negation (0.069 > 0.066) and those that contain a mix of negation and simple statements (0.1427 > 0.1376). There is however, a significant drop in performance, when evaluating on prompts that only contain negation; thus highlighting a major scope of improvement in this regard.
5.5 Improvement over Attention Maps
Finally, inspired from methods such as Attend-and-Excite [5], we visualize attention relevancy maps for both simple and complex spatial prompts and find that we are able to better generate the expected object as well as achieve better spatial localization. In both examples, there are missing objects in the baseline while our model is able to generate them; the bed and the house are not present for Stable Diffusion 2.1 while our model succeeds in generating them. The relevancy map shows that the patches of the image that have high attention to the missing words is spread across the image. We can also see from this figure that the spatial word has high attention to the correct part of the image in our model, compared to the baseline: In our model, as shown in Figure 9 (bottom row), below attends to image patches that are below the bed and right attends to the image patches that are located on the road and on its right, while the Stable diffusion 2.1 model fails to give high relevancy map to the expected locations. We are able to improve upon both the aspects through the intermediate attention maps and the final generated images.
6 Conclusion
In this work, we present findings and techniques that enable improvement of spatial relationships in text-to-image models. We develop a large-scale dataset, SPRIGHT that captures fine-grained spatial relationships across a diverse set of images. Leveraging SPRIGHT, we develop efficient training techniques and achieve state-of-the art performance in generating spatially accurate images. We thoroughly explore various aspects concerning spatial relationships and evaluate the range of diversity introduced by the SPRIGHT dataset. We leave further scaling studies related to spatial consistency as future work. We believe our findings and results facilitate a comprehensive understanding of the interplay between spatial relationships and T2I models, and contribute to the future development of robust vision-language models.
Acknowledgements
We thank Lucain Pouget for helping us in uploading the dataset to the Hugging Face Hub and the Hugging Face team for providing computing resources to host our demo. The authors acknowledge resources and support from the Research Computing facilities at Arizona State University. This work was supported by NSF RI grants #1750082 and #2132724. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
Appendix
Appendix A System Prompt for GPT-4 Evaluation
You are part of a team of bots that evaluates images and their captions. Your job is to come up with a rating between 1 to 10 to evaluate the provided caption for the provided image. Consider the correctness of spatial relationships captured in the provided image. Return the response formatted as a dictionary with two keys: ‘rating’, denoting the numeric rating, and ‘explanation’, denoting a brief justification for the rating.
The captions you are judging are designed to stress-test image captioning programs, and may include:
1. Spatial phrases like above, below, left, right, front, behind, background, foreground (focus most on the correctness of these words).
2. Relative sizes between objects such as small & large, big & tiny (focus on the correctness of these words).
3. Scrambled or misspelled words (the image generator should produce an image associated with the probable meaning). Make a decision as to whether or not the caption is correct, given the image.
A few rules:
1. It is ok if the caption does not explicitly mention each object in the image; as long as the caption is correct in its entirety, it is fine.
2. It is also ok if some captions don’t have spatial relationships; judge them based on their correctness. A caption not containing spatial relationships should not be penalized.
3. You will think out loud about your eventual conclusion. Don’t include your reasoning in the final output.
4. Return the response formatted as a Python-formatted dictionary having two keys: ‘rating’, denoting the numeric rating, and ‘explanation’, denoting a brief justification for the rating.
Appendix B Results on T2I-CompBench
Method Attribute Binding Object Relationship
Color Shape Texture Spatial Non-Spatial
SD 1.4 0.3765 0.3576 0.4156 0.1246 0.3079
SD 2 0.5065 0.4221 0.4922 0.1342 0.3096
Composable v2 [39] 0.4063 0.3299 0.3645 0.0800 0.2980
Structured v2 [14] 0.4990 0.4218 0.4900 0.1386 0.3111
Attn-Exct v2 [5] 0.6400 0.4517 0.5963 0.1455 0.3109
GORS [23] 0.6603 0.4785 0.6287 0.1815 0.3193
DALLE-2 [53] 0.5750 0.5464 0.6374 0.1283 0.3043
SDXL [49] 0.6369 0.5408 0.5637 0.2032 0.3110
PixArt-Alpha [6] 0.6886 0.5582 0.7044 0.2082 0.3179
Kandisnky v2.2 [26] 0.5768 0.4999 0.5760 0.1912 0.3132
DALL-E 3 [45] 0.8110 0.6750 0.8070 - -
Ours (<500 images) 0.6251 0.4648 0.5920 0.2133 0.3132
Table 12: Results on the T2ICompBench Benchmark. a) We achieve state of the art spatial score, across all methods, by efficient fine-tuning on only 444 images. b) Despite not explicitly optimizing for them, we find substantial improvement and competitive performance on attribute binding and non-spatial aspects.
As shown in Table 12, we achieve state of the art performance on the spatial score in the widely accepted T2I-CompBench [23] benchmark. The significance of training on images containing a large number of objects is emphasized by the enhanced performance of our models across various dimensions in T2I-CompBench. Specifically, we enhance attribute binding parameters such as color and texture, alongside maintaining competitive performance in non-spatial aspects.
Appendix C Illustrative Examples from VISOR
In Figure 10, we present illustrative results on the VISOR benchmark.
Appendix D Comparing COCO-30K and Generated Images
In Figure 11, we compare images from COCO, baseline Stable Diffusion and our model. We find that the generated images from our model adhere to the input prompts better, are more photo-realistic in comparison to the baseline.
Appendix E Additional Examples from SPRIGHT
Figure 12 and 13 demonstrate a few correct and incorrect examples present in SPRIGHT. In Figure 13, while most relationships are accurately described in the captions, there are instances where the model struggles to capture the precise spatial nuance.
Appendix F Contributions
AC defined the scope of the project, performed the initial hypothesis experiments and conducted the evaluations. GBMS led all the experimental work and customized the training code. EA generated the dataset, performed the dataset and relevancy map analyses. SP took part in the initial experiments, suggested the idea of re-captioning and performed few of the evaluations and analyses. DG suggested the idea of training with object thresholds and conducted the FAITHScore and GenEval evaluations. TG participated in the early discussions on spatial failures of T2I models. VL, CB, and YZ co-advised the project, initiated and facilitated discussions, and helped shape the the goal of the project. AC and SP wrote the manuscript in consultation with TG, LW, HH, VL, CB, and YZ. All authors discussed the result and provided feedback for the manuscript.
Appendix G Limitations
Since SPRIGHT is a derived dataset, it inherits the limitations of the original datasets. We refer the readers to the respective papers that introduced the original datasets for more details. As shown in our analysis, the generated synthetic captions are not a 100% accurate and could be improved. The improvements can be achieved through better prompting techniques, larger models or by developing methods that better capture low-level image-text grounding. However, the purpose of our work is not to develop the perfect dataset, it is to show the impact of creating such a dataset and its downstream impact in improving vision-language tasks. Since our models are a fine-tuned version of Stable Diffusion, they may also inherit their limitations in terms of biases, inability to generate text in images, errors in generating correct shadow patterns. We present our image fidelity metrics reporting FID on COCO-30K. COCO-30K is not the best dataset to compare against our images, since the average image resolutions in COCO is lesser than those generated by our model which are of dimension 768. Similarly, FID largely varies on image dimensions and has poor sample complexity; hence we also report numbers on the CMMD metric.
Figure 10: Generated images from our model, as described in Section 4.2, on evaluation prompts from the VISOR benchmark.
Figure 11: Illustrative examples comparing ground-truth images from COCO and generated images from Baseline SD 2.1 and our model. The images generated by our model exhibit greater fidelity to the input prompts, while also achieving a higher level of photorealism.
Figure 12: Illustrative examples from the SPRIGHT dataset, where the captions are correct in its entirety; both in capturing the spatial relationships and overall description of the image. The images are taken from CC-12M and Segment Anything.
Figure 13: Illustrative examples from the SPRIGHT dataset, where the captions are not completely correct. The images are taken from CC-12M and Segment Anything.